• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name peter

    Log entry time 08:03:46 on October23,2002

    Entry number 88856

    keyword=shift summary

    00:00 Shift starts with run 2691 in progress; just starting kinematic point Q3_d40 on liquid deuterium. Noticed the 1C12 OTR is out when it ought to be in - correct for the next run.

    04:27 MCC telephoned to say they are having RF problems - the beam is off and they are currently trying to fix the problem.

    04:47 Beam restored.

    05:55 MCC called to say they are having more RF problems - hence we are having more beam trips.

    06:10 Beam restored.

    07:05 MCC called - they took the beam away to try and find the problem causing so many beam trips.

    07:30 MCC claim they are ready to deliver beam again.

    07:40 Beam restored to 80uA; MCC claim this is the highest they can go without tripping the beam at once.

    07:50 Beam current now up to almost 110uA.

    08:00 New shift begins with run # 2699 in progress.

    Run History

    Kinematic Q3_d40

    Run #2691 Good; 1M events. Average current = 91uA
    Run #2692 1C12 OTR now in. Good; 1M; Average current = 94uA
    Run #2693 Good; 1M events. Average current = 93uA
    Run #2694 Good; 1M events. Average current = 83uA
    Run #2695 Good; ended early as beam off; 470k events; average current = 64uA
    Run #2696 Good; 1M events; average current 27uA - lots of trips.
    Run #2697 Short - MCC called with no beam.
    Run #2698 Ended early - lost beam; 762k events.
    Run #2699 Continuing into the new shift.